OpenStudy (anonymous):

1. A sum of money doubles in 20 years on simple interest. It will get triple at the same rate in: a. 40 years b. 50 years c. 30 years d. 60 years e. 80 years

OpenStudy (anonymous):

around 32 years

OpenStudy (anonymous):

if its triple it's going to be 3=(1.03526)^t. Then take the log of each side log3=t*log(1.03526). Therefore, t= log3/log(1.03526)= .4771 /.01505. The answer will be almost 32 years.

OpenStudy (anonymous):

C i think